# Block: Balancing Load in LLM Serving with Context, Knowledge and Predictive Scheduling

### v1

@@ 4.1. Model Instance
A model instance is collection of services deployed on a GPU host that is responsible for model execution and response generation. It consists of two primary components: the main services of the inference framework and a sidecar service called the Predictor. As discussed in §2, different inference frameworks have been widely explored and developed such as (Kwon et al., 2023; Zheng et al., 2024; Hu et al., 2024; Yu et al., 2022; Aminabadi et al., 2022). Similar to Vidur, Block is designed to be framework-agnostic, supporting various frameworks with dynamic batching. Each inference frameworks used by Block can be integrated into it individually by supporting a new status API to export its internal status, such as request lists and GPU memory blocks, for metrics prediction.
The predictor service runs locally on each instance and aggregates runtime data exported by the status API and transforms it into a metrics map via the predict API, consumed by the global scheduler. The Predictor’s main role is to simulate and predict key performance metrics, such as end-to-end latency or TTFT, for incoming requests to be used by Block scheduling. In cases where a request’s actual decoded length exceeds its predicted length, the simulator dynamically adjusts estimated length for prediction by using the monitored decode length plus an additional 10 steps.
The Predictor’s simulator is adapted from Vidur (Agrawal et al., 2024a) in the following way. We redesign the Vidur simulator for single-instance prediction and encapsulated it within the Predictor service. As illustrated in Figure 4, this simulation involves a two-stage process. First, a local scheduler simulator models the batching strategy for a given inference backend. Second, a linear model predicts the execution times for the batches generated in the first stage. Since the model is static and inputs are received on-demand from the status API, the Predictor service is stateless and replicable for single instances to reduce simulation-related overhead through parallelization. Furthermore, the Predictor service is designed with a extendable interface, which could be implemented by alternative simulation frameworks such as SimAI (Wang et al., 2025).
@@ 4.2. Block and Baseline Scheduler
The global scheduler service is designed to be fully distributed and stateless to ensure scalability in large clusters. As illustrated in Figure 4, rather than maintaining a cached, global table of instance statuses, the global scheduler calls the predict API to obtain real-time metrics and predictions for its scheduling decisions. Although this could introduce additional overhead to the end-to-end latency, we anticipate this impact to be minimal, as measured in §6. Besides, the simulation process for scheduling is computationally intensive and relies on data exported from the inference framework. So, we run Predictors locally on instances and leaving global scheduler focus on dispatching only. This approach reduces the overhead associated with both computation and data migration between services. This distributed architecture offers several key advantages. It enhances scheduling efficiency through parallel operations and improves scalability and reliability when serving with large-scale clusters.
@@ 5. Implementation
Block is designed to be backend-agnostic, decoupled from specific backend framework integrations. The current Block prototype works with the integrated vLLM 0.7.2 base version. Integrating a backend involves two main steps. First, the backend’s internal state must be exposed via the new status API as discussed in §4.1. Second, the framework’s local scheduler simulator needs to be implemented to simulate the backend batching strategy. Both steps require minimal effort; for instance, the new vLLM simulator is only 161 lines of code (LoC), and the vLLM API commit is 154 LoC. Besides the integration, to reduce API overhead which could decrease simulation accuracy, we enable vLLM’s multi-process frontend, which separates the inference engine and API frontend into distinct processes. However, parsing dynamic JSON messages between services still incurs constant scheduling overhead (as presented in §6.3), due to Python’s GIL constraints on thread-level parallelism. To further mitigate this overhead, the framework could be migrated to an alternative RPC protocol like gRPC (grp, [n. d.]), or it could leverage Python’s GIL-free features once they become available in FastAPI and vLLM.
@@
  • Min QPM (Queries Per Minute): the default scheduling policy in LiteLLM (lit, [n. d.]), a popular open-source library for routing LLM requests in cloud environments. This policy simply selects instances with minimal latest query-per-minutes.
@@
In addition to the service and simulation functionalities, we also release the training data and scripts for the length estimation model. Consistent with other works with auxiliary models for LLM inference (Gong et al., 2025; Stojkovic et al., 2024; Hu et al., 2024), our model is designed to be lightweight, ensuring its overhead does not significantly impact end-to-end performance. While we initially considered the 7B model from Sequence Scheduling (Zheng et al., 2023b), its training and serving costs were prohibitive to us. We instead fine-tune a RoBERTa-base (Liu et al., 2019) regression model with 125M parameter, which offers a more efficient path to comparable performance. Evaluation is presented in §6.2.
@@
Our current prototype supports key components similarly to other LLM scheduling frameworks. As part of our future work, we plan to add features to align with key trends in LLM serving. While P-D disaggregation has not yet been integrated, we consider its support should be feasible. Similar to LLM-d (git, [n. d.]), it would involve dedicated schedulers for prefill and decode phases to enable inter-phase cache transfer. Given cache transfer is still under active development on inference frameworks, we defer the full implementation and detailed exploration of P-D disaggregation to future work. Additionally, cache transfer enables the exploration of Block Scheduler combined with dynamic rebalancing. Nevertheless, we anticipate that when integrating the above features Block’s performance advantages will persist, as the fundamental scheduling challenges discussed in §3 still remain pertinent in disaggregated settings. Furthermore, incorporating prefix caching (vll, [n. d.]) can enhance simulator accuracy for real-world applications involving multi-turn conversations and repetitive prompts (Zheng et al., 2024). Since it has a negligible impact on experimental setup with de-duplicated dataset, we leave it for further exploration.
@@ 6.1. Experimental Setup
We use ShareGPT (sha, 2023), a real-world dataset consisting of 52K conversations sampled from chatGPT. During our experiments, the prompts in the conversation from shareGPT are sent to the global scheduler following the Poisson distribution under varying arrival rates. We refer this rate as external QPS below.
When serving the model on vLLM, FlashInfer (Ye et al., 2025) is used as the attention library. Also, greedy decoding (Jurafsky and Martin, 2025) is used with temperature set as 0. The maximum request batch size is set to 48, and chunk size for chunked prefilled as 512. This combination shows best performances among all tested configuration sets. To reduce the overhead caused by simulation, each host runs 16 Predictors in parallel to simulate and predict latencies. We observe the such replication reduces the total scheduling overhead up to 50% as described in §6.3.
@@ 6.2.2. Simulation-based Metric Prediction
First, we assess the accuracy of the runtime simulator in predicting metrics and its effectiveness for scheduling, as shown in Figure 5. In this experiment, Block process requests with real length under varying QPS from 20 to 36 using a random scheduler, recording both predicted and actual serving latencies. Online requests have an 1% probability to be sampled. When triggered, a sampled request initiates two actions: 1) reporting the accumulated prediction error for all preceding requests, and 2) broadcasting itself to all instances and marking the instance with minimal predicted latency as selected.
Graphs in the top row show that the prediction error rate is stabilized and is consistently lower when using chunked prefill compared to prioritized prefill. This shows the effectiveness of chuncked prefill in mitigating the disruptive stall bubbles that affect latency prediction. Further, overall error rates between 10% to 15% confirm the simulator’s high accuracy. This 10-15% prediction error is not expected to entirely skew instance selection, as it includes a constant overhead that uniformly impacts all instances and therefore does not alter their relative latency rankings.
@@ 6.3. Request Latency Performance
We use Block* to denote Block operating with predicted lengths from our POC model for latency prediction. As in real-world applications, the actual prompt length could be available by prompt cache (Bang, 2023) with duplicated prompts and estimated lengths are only required for fewer new prompts. To measure single-precision capacity, we conduct further granular search around integer QPS bracketing SLO.
@@
Block* slightly underperforms compared to Block contributed by the error of length estimation but with less overhead addition. This discrepancy is due to the greater uniformity of the estimated output length relative to the actual length, causing higher hit rate on cached batch latencies during simulation. The overhead is below 3% of the e2e latency and tends to decrease once capacity is reached, as e2e latency then spikes more rapidly. Furthermore, INFaaS++ outperforms three basic schedulers under low QPS, but exhibits significant performance degradation with QPS increasing, particularly in tail latencies. Llumnix- mitigates this issue and outperforms other baselines by applying prefill length of pending queries as the correction items over INFaaS++ load calculation as detailed in §5. Taking QPS 32 as an example, Block/Block* reduce average/P99 TTFT by 88.07%/78.6% and 23.58%/10.84% respectively. For e2e latency, Block/Block* achieve reductions of 19.87%/12.56% on mean and 3.55%/0.82% at tails. These improvements ultimately lead to throughput gains of 4.44%/2.53%. Similar trends hold across QPS and get more pronounced at higher QPS.
@@ 6.5. Auto Provisioning
As discussed in §3, provisioning based on runtime metrics could suffer from cold start. To tackle this, predicted metrics can be used to provision instances in a preemptive manner. To this end, we implement two simple auto-provisioning strategies in Block. The preempt strategy adds an instance when predicted latency reaches 70s, while relief strategy means provisioning only when actual latency hits the same threshold. We conduct experiments starting with six available instances under a QPS of 24. We also establish a baseline with a sufficient static cluster of 10 instances.
Results in Figure 8(a) support our hypothesis that a relief strategy causes over-provisioning: as the newly added hosts are unable to relief queued requests, which could trigger further provisioning, all backup instances are quickly exhausted, and this causes memory imbalances and GPU resource wastage. In contrast, preemptive provisioning activates earlier than in the case of a relief approach, leading to smoother changes in cluster size and ultimately provisioning only the necessary instances. This results in lower variance, higher GPU utilization, 20.1% reduction in P99 latency (89.2/71.2) and 81% requests over 70 seconds (627/129).
@@ Appendix A Latency CDF for selected QPS
We present the CDF distribution for TTFT and end-to-end latency for selected QPS as 20, 24 and so in Figure 9, which clearly confirm Block/Block* advantage on reduction tail latencies with high QPS.

### v2

@@ 4.1. Model Instance
A model instance is collection of services deployed on a GPU host that is responsible for model execution and response generation. It consists of two primary components: the main services of the inference framework and a sidecar service called the Predictor. As discussed in §2, different inference frameworks have been widely explored and developed such as (Kwon et al., 2023; Zheng et al., 2024; Hu et al., 2024; Yu et al., 2022; Aminabadi et al., 2022). Similar to Vidur, Block is designed to be a framework-agnostic, supporting various frameworks with dynamic batching. Each inference framework used by Block can be integrated into it individually by supporting a new status API to export its internal status, such as request lists and GPU memory blocks, for metrics prediction.
The predictor service runs locally on each instance and aggregates runtime data exported by the status API and transforms it into a metrics map via the predict API, consumed by the global scheduler. The Predictor’s main role is to simulate and predict key performance metrics, such as end-to-end latency or TTFT, for incoming requests to be used by Block scheduling. In cases where a request’s actual decoded length exceeds its predicted length, the simulator dynamically adjusts the estimated length for prediction by using the monitored decode length plus an another 10 steps.
The Predictor’s simulator is adapted from Vidur (Agrawal et al., 2024a) in the following ways. We redesign the Vidur simulator for single-instance prediction and encapsulated it within the Predictor service. As illustrated in Figure 4, this simulation involves a two-stage process. First, a local scheduler simulator models the batching strategy for a given inference backend. Second, a linear model predicts the execution times for the batches generated in the first stage. Since the model is static and inputs are received on-demand from the status API, the Predictor service is stateless and replicable for single instances to reduce simulation-related overhead through parallelization. Furthermore, the Predictor service is designed with an extendable interface, which could be implemented by alternative simulation frameworks such as SimAI (Wang et al., 2025).
@@ 4.2. Block and Baseline Scheduler
The global scheduler service is designed to be fully distributed and stateless to ensure scalability in large clusters. As illustrated in Figure 4, rather than maintaining a cached, global table of instance statuses, the global scheduler calls the predict API to obtain real-time metrics and predictions for its scheduling decisions. Although this could introduce additional overhead into the end-to-end latency, we anticipate this impact to be minimal, as measured in §6. Besides, the simulation process for scheduling is computationally intensive and relies on data exported from the inference framework. So, we run Predictors locally on instances and leave global scheduler focus on dispatching only. This approach reduces the overhead associated with both computation and data migration between services. This distributed architecture offers several key advantages. It enhances scheduling efficiency through parallel operations and improves scalability and reliability when serving with large-scale clusters.
@@ 5. Implementation
Block is designed to be backend-agnostic, decoupled from specific backend framework integration. The current Block prototype works with the integrated vLLM 0.7.2 base version. Integrating a backend involves two main steps. First, the backend’s internal state must be exposed via the new status API as discussed in §4.1. Second, the framework’s local scheduler simulator needs to be implemented to simulate the backend batching strategy. Both steps require minimal effort; for instance, the new vLLM simulator is only 161 lines of code (LoC), and the vLLM API commit is 154 LoC. Besides the integration, to reduce API overhead which could decrease simulation accuracy, we enable vLLM’s multi-process frontend, which separates the inference engine and API frontend into distinct processes. However, parsing dynamic JSON messages between services still incurs constant scheduling overhead (as presented in §6.3), due to Python’s GIL constraints on thread-level parallelism. To further mitigate this overhead, the framework could be migrated to an alternative RPC protocol like gRPC (grp, [n. d.]), or it could leverage Python’s GIL-free features once they become available in FastAPI and vLLM.
@@
  • Min QPM (Queries Per Minute): the default scheduling policy in LiteLLM (lit, [n. d.]), a popular open-source library for routing LLM requests in cloud environments. This policy simply selects instances with minimal latest QPM.
@@
In addition to the service and simulation functionalities, we also release the training data and scripts for the length estimation model. Consistent with other works with auxiliary models for LLM inference (Gong et al., 2025; Stojkovic et al., 2024; Hu et al., 2024), our model is designed to be lightweight, ensuring its overhead does not significantly impact end-to-end performance. While we initially considered the 7B model from Sequence Scheduling (Zheng et al., 2023b), its training and serving costs were prohibitive for us. We instead fine-tune a RoBERTa-base (Liu et al., 2019) regression model with 125M parameter, which offers a more efficient path to comparable performance. Evaluation is presented in §6.2.
@@
Our current prototype supports key components similarly to other LLM scheduling frameworks. As part of our future work, we plan to add features to align with key trends in LLM serving. While P-D disaggregation has not yet been integrated, we consider its support should be feasible. Similar to LLM-d (git, [n. d.]), it would involve dedicated schedulers for prefill and decode phases to enable inter-phase cache transfer. Given cache transfer is still under active development on inference frameworks, we defer the full implementation and detailed exploration of P-D disaggregation to future work. Additionally, cache transfer enables the exploration of Block Scheduler combined with dynamic rebalancing. Nevertheless, we anticipate that when integrating the above features Block’s performance advantages will persist, as the fundamental scheduling challenges discussed in §3 remain pertinent in disaggregated settings. Furthermore, incorporating prefix caching (vll, [n. d.]) can enhance simulator accuracy for real-world applications involving multi-turn conversations and repetitive prompts (Zheng et al., 2024). Since it has a negligible impact on experimental setup with de-duplicated dataset, we leave it for further exploration.
@@ 6.1. Experimental Setup
We use ShareGPT (sha, 2023), a real-world dataset consisting of 52K conversations sampled from chatGPT. During our experiments, the prompts in the conversation from ShareGPT are sent to the global scheduler in order following the Poisson distribution under varying arrival rates. We refer this rate as external QPS below.
When serving the model on vLLM, FlashInfer (Ye et al., 2025) is used as the attention library. Also, greedy decoding (Jurafsky and Martin, 2025) is used with temperature set as 0. The maximum request batch size is set to 48, and chunk size for Chunked Prefill as 512. This combination shows best performances among all tested configuration sets. To reduce the overhead caused by simulation, each host runs 16 Predictors in parallel to simulate and predict latencies. We observe such replication reduces total scheduling overhead up to 50% as described in §6.3.
@@ 6.2.2. Simulation-based Metric Prediction
First, we assess the accuracy of the runtime simulator in predicting metrics and its effectiveness for scheduling, as shown in Figure 5. In this experiment, Block process requests with real length under varying QPS from 20 to 36 using a random scheduler, recording both predicted and actual serving latencies. Online requests have a 1% probability to be sampled. When triggered, a sampled request initiates two actions: 1) reporting the accumulated prediction error for all preceding requests, and 2) broadcasting itself to all instances and marking the instance with minimal predicted latency as selected.
Graphs in the top row show that the prediction error rate is stabilized and is consistently lower when using Chunked Prefill compared to prioritized prefill. This shows the effectiveness of Chunked Prefill in mitigating the disruptive stall bubbles that affect latency prediction. Further, overall error rates between 10% to 15% confirm the simulator’s high accuracy. This 10-15% prediction error is not expected to entirely skew instance selection, as it includes a constant overhead that uniformly impacts all instances and therefore does not alter their relative latency rankings.
@@ 6.3. Request Latency Performance
We use Block* to denote Block operating with predicted lengths from our POC model for latency prediction. As in real-world applications, the actual prompt length could be available by prompt cache (Bang, 2023), with duplicated prompts and estimated lengths are only required for fewer new prompts. To measure single-precision capacity, we conduct further granular search around integer QPS bracketing SLO.
@@
Block* slightly underperforms compared to Block due to the error of length estimation but with less overhead addition. This discrepancy is due to the greater uniformity of the estimated output length relative to the actual length, causing higher hit rate on cached batch latencies during simulation. The overhead is below 3% of the e2e latency and tends to decrease once capacity is reached, as e2e latency then spikes more rapidly. Furthermore, INFaaS++ outperforms three basic schedulers under low QPS, but exhibits significant performance degradation with QPS increasing, particularly in tail latencies. Llumnix- mitigates this issue and outperforms other baselines by applying prefill length of pending queries as correction items over INFaaS++ load calculation as detailed in §5. Taking QPS 32 as an example, Block/Block* reduce average/P99 TTFT by 88.07%/78.6% and 23.58%/10.84% respectively. For e2e latency, Block/Block* achieve reductions of 19.87%/12.56% on mean and 3.55%/0.82% at tails. These improvements ultimately lead to throughput gains of 4.44%/2.53%. Similar trends hold across QPS and get more pronounced at higher QPS.
@@ 6.5. Auto Provisioning
As discussed in §3, provisioning based on runtime metrics could suffer from cold start. To tackle this, predicted metrics can be used to provision instances in a preemptive manner. To this end, we implement two simple auto-provisioning strategies in Block. The preempt strategy adds an instance when predicted latency reaches 70 seconds, while relief strategy means provisioning only when actual latency hits the same threshold. We conduct experiments starting with six available instances under a QPS of 24. We also establish a baseline with a sufficient static cluster of 10 instances.
Results in Figure 8(a) support our hypothesis that a relief strategy causes over-provisioning: as the newly added hosts are unable to relief queued requests, which could trigger further provisioning, all backup instances are quickly exhausted, and this causes memory imbalances and GPU resource wastage. The preemptive strategy activates earlier and lead to smoother changes in cluster size and ultimately provisioning only the necessary instances. This results in lower variance and higher utilization for memory overtime and reducing P99 latency by 20.1% (89.2/71.2) and requests over 70 seconds by 81% (627/129) with less instances.
@@ Appendix A Latency CDF for selected QPS
We present the CDF distribution for TTFT and end-to-end latency for selected QPS as 20, 24 and so on in Figure 9, which clearly confirm Block/Block* advantage on reduction tail latencies with high QPS.